OpenStudy (anonymous):

tony's club is selling oranges to raise money. for every box they sell, they get 1 1/8 dollars profit. They have sold 75 boxes already. how many more boxes must they sell to raise 180 dollars ?

OpenStudy (texaschic101):

1 1/8 x 75 = 9/8 x 75 = 675/8 = 84.37 round down to 84 Tony has already made $ 84 from selling 75 boxes. Now set up a proportion... 75/84 = x/180 (75 boxes made $84 = x boxes made $180 cross multiply... 84x = (75)(180) 84x = 13500 x = 13500/84 x = 160.7 = 160 boxes check... 160 x 1 1/8 = 160 x 9/8 = 1440/8 = 180 ANSWER : Tony would have to sell 160 more boxes,

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

i just had the same question and that was the wrong answer

OpenStudy (anonymous):

it is 85 boxes total

OpenStudy (anonymous):

i had the same question and that is the right answer